# Walking the Tightrope of LLMs for Software Development: A Practitioners’ Perspective

### v1

@@ 4.2 Directions for Future Work
Based on the findings from our analysis, we have identified research gaps involved to software practitioners adopting LLMs for Software Development, not explored in our current work.
@@
How do LLMs empower developers’ time? Participants mentioned different ways LLMs empower their time. While some participants would just proceed to the next task after finish current tasks using LLMs, one of the participants would use LLMs to give an extra time before close tasks, e.g.: 9 “The big advantage is time. You can even bargain for a bit of time too. For example, suppose I have a task that I know will take a day to do, if the AI can do it for me, I won’t deliver it quickly. I’ll take my time and at the end of the day I’ll hand it in.” – P4 [Software Developer].
How does laziness drive developers’ interaction with LLMs? We identified some examples that point to laziness in software developers’ interaction with LLMs. Software developers might feel lazy to write prompts, and interaction with LLMs can also make software developers lazy.
@@
How do LLMs affect developers’ intuition? According to Naur [54], “Software development in all its phases, and irrespective of the techniques employed in its pursuits must and will always depend on intuition”. Intuition built from previous experiences acts as a safeguard mechanism, enabling software developers to navigate LLM’s wrong suggestions. While Chen et al. [55] investigated the role of human intuition when using LLMs, the study regarding the effects of LLMs on human intuition remains unexplored.
How do LLMs contribute to burnout in software projects? Investigations into burnout in SE have been the focus of numerous studies [56]. In their systematic mapping study, Tulili et al. [56] identified the following causes for burnout: personality traits (e.g., neuroticism), work-related factors (e.g., job demand, job overload), communication practices, agile practices, and physiological factors. In our study, one of our participants expressed concern about organisations pressuring developers to expand their responsibilities by using LLMs as justification, which may contribute to developer burnout.
@@
Controversial use cases. Participants demonstrate different opinions regarding using LLMs for test generation, code generation, and debugging. LLMs are discouraged for test-related tasks due to providing the illusion of false assurance. Wang et al. [7] argue that using LLMs for integration tests may result in errors or unreliable results because those tests may exceed the capacity of the LLM to process and analyse. LLMs are also discouraged for code generation due to the harmful long-term effects towards software developers’ skills. Some participants also highlight LLMs may provide general debugging guidance, since they cannot run the code.
@@ 5 Related Work
Figure 5: Comparison with related work focused on benefits and disadvantages. The colored bars represent the overlap with existing literature [57, 58, 59, 60], and empty bars indicate new findings.
@@
Banh et al. [58] conducted eighteen interviews with IT-related professionals (i.e., software developers, a product owner, and a scrum master) between August 2023 and January 2024. They aimed to understand how to integrate Generative AI in Software Engineering, exploring the opportunities and challenges related to adopting Generative AI tools. They applied the traditional Grounded Theory, resulting in the conceptual framework of generative AI integration in Software Engineering practices. Their findings, similar to ours, are: reduced development time, issues with intellectual property, underestimated overhead, developer empowerment, and code quality improvement via LLMs. However, our findings also include the contractions involving the benefits and disadvantages, such as code degradation via LLMs, LLMs dragging tasks (see 3.2.1). Our results point out the necessity for a balanced approach to the adoption of LLMs, advocating for their use as a net positive in the software development process. Our recommendations in Section 3.3 contain practical suggestions contributing to achieve this.
Liang et al. [59] employed traditional Grounded Theory [62] through twenty interviews and observation of fifteen software developers. We could not identify when interviews and observation were conducted. Their contribution includes an understanding of prompt programming practices. Similar to our study, they identified the potential of LLMs to assist developers in improving their mental model. Although their research focused on prompt programming aspects, they present their impact across different software development tasks similar to us (See Table 3.1.1 and 3.2.1). Our findings distinguish by showing not only the impact of software practitioners’ interaction with LLMs, but also how it reflects in others levels, such as team level (e.g., mitigating interruptions).
@@ 6 Evaluation
6.1 Evaluating STGT Application
The evaluation of the STGT method application consists of credibility and rigour as key criteria [25, 27]. In terms of credibility, our section 2 provides details involving the recruitment process (social media, and emailing), sampling method (purposive sampling followed by convenient sampling), how iterative and interleaved data occurred (three rounds of data collection and analysis), and how memos was written and applied (to guide structure of emerging concepts and sub-categories, and as directions for future work). Concerning rigour, our section 2 also provides examples of basic coding and constant comparison (See Fig. 2) and embedded sanitised evidence (i.e., several interview quotes throughout Section 3).
6.2 Evaluating STGT Outcomes
Findings from the application of STGT for data analysis should exhibit originality, relevance, and density [25, 27]. In terms of originality, we position our findings within comparison to related work in Section 5. The relevance of understanding the impact of LLMs on software practitioners is highlighted in Section 1. Regarding density, Section 3 condenses the sub-categories and concepts that compose the category “Impact on using LLMs”. We illustrate the different aspects by using several interview quotes.
@@ 7 Threats to Validity
We will discuss the study limitations using the Total Quality Framework (TQF) developed by [63], suitable for qualitative studies in software engineering [64, 65]. This framework is structured in the following aspects: credibility regarding data collection, analyzability on the data, transparency of reporting, and usefulness of the findings.
@@
Usefulness. It refers to how useful the study contributions and findings are. Our findings can support software team leaders and IT managers in evaluating whether LLMs align with their context and needs, as well as guiding SE researchers for future research needs.
@@ 8 Conclusion
In this paper, we present the software practitioners’ perspective on the impact of LLMs at the individual, team, organisation, and society levels. We also present suggestions on how to balance the benefits and disadvantages of using LLMs. Throughout the paper, we present LLMs as a double-edged sword. At the individual level, most of the participants mention LLMs boosting software development, saving time, and reducing effort as benefits; and LLMs slowing software development, increasing effort as a downside. At the team level, most of the participants refer to LLMs as mitigating interruptions, but also reducing mentorship opportunities. At the organisation level, participants highlight the cost savings due to LLMs, but also security and privacy concerns. At the society level, participants mention LLMs promoting entrepreneurship, but also the erosion of social trust. We conclude that balanced control is the ideal approach towards LLMs. Our findings also indicate LLMs are improving developer experience, enabling them to reach flow state.
@@ 9 Acknowledgment
Samuel is supported by a Faculty of IT PhD scholarship. We would like to express profound gratitude to all participants who took part in this study, Hashini Gunatilake, Humphrey Obie, and Marcos Medeiros for the wonderful feedback on this research, and Nimmi Weeraddana and Haoyu Gao for helping during participant recruitment.

### v2

@@ 4.2 Implication for Research
    Based on the findings, we identified research gaps related to software practitioners adopting LLMs for Software Development.
    What factors influence developers’ decisions to delegate SE tasks to LLMs? Our interview participants described adopting different attitudes regarding the use of LLMs for SE tasks. We observed that task risk and developers’ motivation play an undeniable role in task delegation to LLMs. For instance, Participant P19 argues against relying on LLMs for test generation, warning that it could create a false sense of security. Regarding developers’ motivation, a lack of enthusiasm for certain tasks (e.g., updating documentation) among developers seems to encourage them to delegate such responsibilities. We suggest further studies to investigate other factors and how they interplay with task risk and developers’ motivation.
    How do LLMs empower developers during tight deadlines? Our data reveals a spectrum of behaviours: whereas some developers use LLMs to meet tight deadlines rapidly, others may inflate their effort estimations when motivation is low. Researchers should further examine the triggers for such inflation to better understand the long-term impact of AI assistance on software process metrics.
@@
    How do LLMs affect developers’ intuition? According to Naur [44], “Software development in all its phases, and irrespective of the techniques employed in its pursuits, must and will always depend on intuition”. Intuition built from previous experiences acts as a safeguard mechanism, enabling software developers to navigate LLMs’ wrong suggestions. While Chen et al. [7] investigated the role of human intuition when using LLMs, the study regarding the effects of LLMs on human intuition remains unexplored.
    How does agentic software engineering contribute to developers’ burnout? Investigations into burnout in SE have been the focus of numerous studies [56]. In their systematic mapping study, Tulili et al. [56] identified the following causes for burnout: personality traits (e.g., neuroticism), work-related factors (e.g., job demand, job overload), communication practices, agile practices, and physiological factors. However, software developers’ burnout from adopting LLM tools is an emergent topic that requires attention from researchers. A recent large survey conducted by Feng et al. [17] found that job demands—such as organisational pressures and increased workload—are positively associated with burnout, while job resources related to generative AI adoption—such as autonomy and access to learning resources—are negatively associated with it. Beyond the adoption of generative AI, the rise of agentic software development can place a considerable burden on developers, who may need to supervise several agentic tools doing different software engineering tasks in parallel. Future research should investigate how this expanding supervisory burden interacts with established burnout drivers — and critically, what constitutes a sustainable threshold of concurrent agent supervision for software developers.
    How does agentic software engineering influence job title inflation within the tech industry? Title inflation occurs when employees are granted senior-level designations that exceed their actual duties or compensation [48]. In the era of agentic workflows, developers across all seniority levels are increasingly tasked with overseeing autonomous agents in addition to their traditional responsibilities. However, as these roles evolve to include agent orchestration, compensation structures often remain static. We propose that future studies investigate whether this gap between responsibility and reward is being bridged by inflated job titles and what impact this has on the long-term transparency of the software labour market.
    What are the most appropriate LLM maturity model stages for organisations? Our data shows varying degrees of LLM adoption, ranging from superficial tool access (e.g., GitHub Copilot for code completion) to deep integration within DevOps pipelines and code review processes. While the Quotient maturity framework [47] identifies five stages—ranging from ad hoc adoption to end-to-end autonomy—and suggests that stages 3 (standardised workflows) and 4 (supervised automation) are the most appropriate targets, these recommendations lack broad empirical support. We advocate for longitudinal studies to validate which maturity levels actually yield the highest return on investment and team stability in practice.
@@
    Controversial use cases. Participants demonstrate different opinions regarding using LLMs for test generation, code generation, and debugging. LLMs are discouraged for test-related tasks due to providing the illusion of false assurance. Wang et al. [57] argue that using LLMs for integration tests may result in errors or unreliable results because those tests may exceed the capacity of the LLM to process and analyse. LLMs are also discouraged for code generation due to the harmful long-term effects towards software developers’ skills. Some participants also highlight that LLMs may provide general debugging guidance, since they cannot run the code.
@@ 5 Related Work
    Figure 5: Comparison with related work focused on benefits and challenges. The colored bars represent the overlap with existing literature [5, 4, 35, 34], and empty bars indicate new findings.
@@
    Banh et al. [4] conducted eighteen interviews with IT-related professionals (i.e., software developers, a product owner, and a scrum master) between August 2023 and January 2024. They aimed to understand how to integrate Generative AI in Software Engineering, exploring the opportunities and challenges related to adopting Generative AI tools. They applied the traditional Grounded Theory, resulting in the conceptual framework of generative AI integration in Software Engineering practices. Their findings, similar to ours, are: reduced development time, issues with intellectual property, underestimated overhead, developer empowerment, and code quality improvement via LLMs. However, our findings also include the contradictions involving the benefits and challenges, such as code degradation via LLMs, LLMs dragging tasks (see LABEL:subsubsec:disruptflow). Our results point out the necessity for a balanced approach to the adoption of LLMs, advocating for their use as a net positive in the software development process. Our recommendations in Section LABEL:subsec:recommendations contain practical suggestions contributing to achieve this.
    Liang et al. [35] employed traditional Grounded Theory [26] through twenty interviews and observation of fifteen software developers. We could not identify when interviews and observation were conducted. Their contribution includes an understanding of prompt programming practices. Similar to our study, they identified the potential of LLMs to assist developers in improving their mental model. Although their research focused on prompt programming aspects, they present their impact across different software development tasks similar to us (See Table 3.1.1 and LABEL:tab:disadvantageIndividualLevel). Our findings distinguish by showing not only the impact of software practitioners’ interaction with LLMs, but also how it reflects in others levels, such as team level (e.g., mitigating interruptions).
@@ 6 Evaluation
    6.1 Evaluating STGT4DA Application
    The evaluation of the STGT4DA application consists of credibility and rigour as key criteria [22, 23]. In terms of credibility, our section 2 provides details involving the recruitment process (social media and emailing), sampling method (purposive sampling followed by convenient sampling), how iterative and interleaved data occurred (three rounds of data collection and analysis), and how memos were written and applied (to guide the structure of emerging concepts and sub-categories, and as directions for future work). Concerning rigour, our section 2 also provides examples of basic coding and constant comparison (See Fig. 2) and embedded sanitised evidence (i.e., several interview quotes throughout Section 3).
    6.2 Evaluating STGT4DA Outcomes
    Findings from the application of STGT4DA should exhibit originality, relevance, and density [22, 23]. In terms of originality, we position our findings in comparison to related work in Section 5. The relevance of understanding the impact of LLMs on software practitioners is highlighted in Sections 1, 4.1, and 4.2. Regarding density, Section 3 condenses the sub-categories and concepts that compose the category “Impact on using LLMs”. We illustrate the depth of the findings through multiple levels of codes, concepts, and sub-categories, and the richness of findings through various aspects of the phenomenon, illustrated through numerous interview quotes.
@@ 7 Threats to Validity
    We will discuss the study limitations using the Total Quality Framework (TQF) developed by [49], suitable for qualitative studies in software engineering [33, 21]. This framework is structured in the following aspects: credibility regarding data collection, analyzability of the data, transparency of reporting, and usefulness of the findings.
@@
    Usefulness. It refers to how useful the study’s contributions and findings are. Our findings can support software team leaders and IT managers in evaluating whether LLMs align with their context and needs, as well as guiding SE researchers for future research needs. We acknowledge that the fast-paced evolution of LLMs can quickly make some of our findings outdated. However, empirical studies help capture and document the evolving state of practice, especially as previous models become inaccessible. Further investigations can compare to our findings and report on the latest LLM capabilities.
@@ 8 Conclusion
    In this paper, we present the software practitioners’ perspective on the impact of LLMs at the individual, team, organisation, and society levels. Our findings comprehend benefits and challenges of using LLMs for software development. At the individual level, most of the participants mention LLMs saving time and reducing effort as benefits, and frictions and overhead introduced by LLM-assisted software development as challenges. At the team level, most of the participants refer to LLMs as mitigating interruptions, but also reducing mentorship opportunities. At the organisation level, participants highlight the cost savings due to LLMs, but also security and privacy concerns. At the society level, participants mention LLMs promoting entrepreneurship, but also the erosion of social trust. The challenges identified create an imbalance in the use of LLMs. To address this issue, our findings also include actionable suggestions to help mitigate this imbalance and support developers as they navigate the integration of LLMs into software development.
@@ 9 Acknowledgment
    Samuel is supported by a Faculty of IT PhD scholarship. We would like to express profound gratitude to all participants who took part in this study, Hashini Gunatilake, Humphrey Obie, and Marcos Medeiros, for the wonderful feedback on this research, and Nimmi Weeraddana and Haoyu Gao for helping during participant recruitment.